CIIT EXAMINATION 2010

Name ( )

1. In Traveling Salesperson Problem (TSP) with 5 cities A, B, C, D and E, what is a possible interpretation of the chromosome (1, 4, 3, 5, 2) such that any crossover will create a legitimate tour, and why you think so?

2. In Iterated Prisonner’s Dillemma (IPD), assuming your chromosome is

(1111000011110000111100001111000011110000111100001111000011110000)

and history of the game was 1-0-1-0-1-0. (1) What should be your next action? And why? (2) Then the oponent’s next two action was 1 and 0, what is your best next twp actions and why?

3. To solve Sorting Network Problem with 8 items, you have a chromosome (3 7 2 4 8 5 6 1). (1) Give your interpretation of this chromosome and, (2) Sort 8 integers 2 3 4 8 5 7 1 6 from small to large.

4. In evolving Finite State Machine by GA, why the number of states should be like 2,4,8,16,32,... rather than 7 for example?

5. Try to propose your very original problem in which you can apply GA (evolving chromosomes). (Write it behind this paper as presise as possible.)